Standard Direct Materials Cost per Unit

Roanoke Company produces chocolate bars. The primary materials used in producing chocolate bars are cocoa, sugar, and milk. The standard costs for a batch of chocolate (1,850 bars) are as follows:

Ingredient Quantity Price
Cocoa 540 lbs. $0.30 per lb.
Sugar 150 lbs. $0.60 per lb.
Milk 120 gal. $1.60 per gal.

Determine the standard direct materials cost per bar of chocolate. If required, round to the nearest cent.
$per bar
Answer #1

Answer:-

Total direct material cost of a batch of 1850 bars

Cocoa . 540 lbs*$0.30= $162

Sugar= 150 lbs*$0.60=$ 90

Milk= 120 gal*$1.60=$192

Total cost of a batch of 1850 bars= 162+90+192= $444

Direct material cost per bar = $444 ÷ 1850= 0.24 $ or 24 cents

The answer is 0.24 $ or 24 cents.

if you have any query please ask me in comment box i am here to helps you don't give direct Thumbs down.if you satisfied my work give Thumbs UP

                                                *******THANK YOU********